# A Clique Partitioning-Based Algorithm for Graph Compression

### v2

@@ Abstract
Reducing the running time of graph algorithms is vital for tackling real-world problems such as shortest paths and matching in large-scale graphs, where path information plays a crucial role. This paper addresses this critical challenge of reducing the running time of graph algorithms by proposing a new graph compression algorithm that partitions the graph into bipartite cliques and uses the partition to obtain a compressed graph having a smaller number of edges while preserving the path information. This compressed graph can then be used as input to other graph algorithms for which path information is essential, leading to a significant reduction of their running time, especially for large, dense graphs. The running time of the proposed algorithm is $O(mn^{\delta})$, where $0\leq\delta\leq 1$, which is better than $O(mn^{\delta}\log^{2}n)$, the running time of the best existing clique partitioning-based graph compression algorithm (the Feder-Motwani (FM) algorithm). Our extensive experimental analysis show that our algorithm achieves a compression ratio of up to $26\%$ greater and executes up to 105.18 times faster than the FM algorithm. In addition, on large graphs with up to 1.05 billion edges, it achieves a compression ratio of up to 3.9, reducing the number of edges up to $74.36\%$. Finally, our tests with a matching algorithm on sufficiently large, dense graphs, demonstrate a reduction in the running time of up to 72.83% when the input is the compressed graph obtained by our algorithm, compared to the case where the input is the original uncompressed graph.
@@ 1 Introduction
Graphs are versatile and powerful when used to model and analyze complex real-world problems in various domains, such as social, biological, and communication networks. As the volume of data generated and stored continues to grow, graphs become increasingly large and complex, presenting significant challenges for processing and analyzing data, including prohibitively high processing times and huge storage requirements. To address these challenges, graph compression techniques such as edge compression are employed to decrease the size of a graph while maintaining its essential properties. Edge compression involves shrinking the edge set of a graph, resulting in a compressed graph with fewer edges than the original. During edge compression it is important to preserve the path information that plays a crucial role in graph’s connectivity, i.e., the ability to reach one vertex from another using a sequence of edges. If the path information is preserved during compression, the compressed graphs can be used as input to several other graph algorithms for which the path information is essential such as matching and all-pairs shortest path algorithms, leading to a significant reduction of their running time, especially for large, dense graphs.
@@
However, to benefit from using the compressed graph as input to a graph algorithm, it is critical that the compression algorithm and the obtained compressed graph meet the following requirements: 1) the compression algorithm must have a low running time so that when used as a preprocessing step for other graph algorithms, like matching, leads to a lower overall execution time than in the case of executing the graph algorithms on the original graph; and 2) the compressed graphs are directly usable as input to other graph algorithms or require minimal modifications by these algorithms. These conditions make the design of graph compression algorithms challenging. In this paper, we address this challenge by focusing on bipartite graphs and propose a new Clique Partitioning-based Graph Compression (CPGC) algorithm that compresses bipartite graphs while preserving their path information. A bipartite graph is a graph where the vertices can be divided into two disjoint sets, such that no two vertices within the same set are adjacent. We theoretically and experimentally show that our algorithm improves the running time and the compression rate over the algorithm proposed by Feder and Motwani [9], i.e., the best existing clique partitioning-based graph compression algorithm used to speed-up the execution of other graph algorithms.
1.1 Related Work.
The problem of representing a graph in a compressed form has been extensively explored, with techniques typically falling into two categories: lossless and lossy compression [2]. The lossless compression retains the entirety of graph information, while the lossy compression sacrifices some details during the compression process. As one of the key requirements of our approach is to preserve the path information, it falls into the lossless compression techniques category. The majority of the proposed lossless graph compression techniques focus on specific domains such as web graphs, social networks, and chemistry networks [2]. Those techniques primarily focus on reducing the storage space [4] and/or efficiently addressing specific graph queries, such as graph pattern matching, community query, etc. [1, 3, 5, 8, 13]. Buehrer and Chellapilla [3] achieved high compression ratios and small depths by using bipartite cliques with frequent itemset mining for compressing web graphs. However, while these techniques achieve high compression ratios, the resulting compressed graphs may not be directly usable as input to algorithms like all-pairs shortest paths, often requiring significant modifications or reconstruction overhead. Similarly, Fan et al. [8] achieved significant compression of large graphs by converting stars, cliques, and paths into super-nodes, achieving speedups in query tasks like shortest distance computations. While this technique is compatible with algorithms like matching, obtaining the path information for all nodes may introduce high overhead, increasing the total running time. While compressed graphs tailored for specific query tasks demonstrate efficiency, seamlessly integrating them with certain graph algorithms, such as matching, remains a challenge due to the critical need for comprehensive path information.
To the best of our knowledge the closest work to ours is that of Feder and Motwani [9] who proposed a graph compression technique that preserves the path information of the original graph while retaining the algorithmic properties of the graph. Using the compressed graph obtained by their technique as input to other graph algorithms, such as bipartite matching [11], edge connectivity [12], and vertex connectivity [7], leads to a significant reduction in their running time. The technique of Feder and Motwani [9] is based on partitioning the graph into bipartite cliques (complete bipartite subgraphs), that is, a set of bipartite cliques that partition the edge set of the graph. They provided an algorithm for compressing bipartite graphs based on partitioning the graph into bipartite cliques that runs in $O(mn^{\delta}\log^{2}n)$ time, where $n$ is the number of vertices, $m$ is the number of edges, and $\delta$ is a constant such that $0\leq\delta\leq 1$. In the following section, we present Feder and Motwani’s work as the background for clique partitioning-based graph compression and provide the motivation for our approach.
@@ 2 Background and Motivation
2.1 Background.
Feder and Motwani [9] proposed a graph compression algorithm (called FM in the rest of the paper) that involves finding $\delta$-cliques and replacing each of them with a tripartite graph having two of the partitions the same as those of the clique it replaces, and the third partition composed of a single newly added vertex. A $\delta$-clique in a bipartite graph graph $G(U,W,E)$ with $|U|=|W|=n$ and $|E|=m$, is a complete bipartite subgraph with the left partition $U^{\prime}$ of size $\lceil n^{1-\delta}\rceil$ and the right partition $W^{\prime}$ of size ${k}(n,m,\delta)=\Big{\lfloor}{\frac{\delta\log n}{\log(2n^{2}/{m})}}\Big{\rfloor}$, where $\delta$ is a constant such that $0\leq\delta\leq 1$. A tripartite graph $G^{*}=(U,W,Z,E^{*})$ consists of three disjoint sets of vertices $U$, $W$, and $Z$, $(U\cap W=W\cap Z=U\cap Z=\emptyset)$, and a set of edges $E^{*}$, where each edge connects either a vertex in $U$ to a vertex in $W$, or a vertex in $W$ to a vertex in $Z$, or a vertex in $U$ to a vertex in $Z$. Figure 1(a) and 1(c) show an example of a $\delta$-clique with left partition $\{u_{1},u_{2},u_{3},u_{4},u_{5},u_{7},u_{8}\}$ and right partition $\{w_{4},w_{5}\}$ in a bipartite graph $G$, and the corresponding tripartite graph that replaces it in the compressed graph $G^{*}$, respectively. The number of edges in the $\delta$-clique is $14=7\times 2$ while the number of edges in the corresponding tripartite graph is $9=7+2$, thus the number of edges in the compressed graph is reduced by 5.
$u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{5}$$u_{6}$$u_{7}$$u_{8}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$$w_{7}$$w_{8}$

### v3

# Speeding-up Graph Algorithms via Clique Partitioning

@@ Abstract
Reducing the running time of graph algorithms is vital for tackling real-world problems such as shortest paths and matching in large-scale graphs, where path information plays a crucial role. To address this critical challenge, this paper introduces a graph restructuring algorithm that identifies bipartite cliques and replaces them with tripartite graphs. This restructuring leads to fewer edges while preserving complete graph path information, enabling the direct application of algorithms like matching and all-pairs shortest paths to achieve significant runtime reductions, especially for large, dense graphs. The running time of the proposed algorithm for a graph $G(V,E)$, with $|V|=n$ and $|E|=m$ is $O(mn^{\delta})$, which is better than $O(mn^{\delta}\log^{2}n)$, the running time of the best existing algorithm for speeding-up other graph algorithms (the Feder-Motwani (FM) algorithm), where $0\leq\delta\leq 1$. Both the FM algorithm and the proposed algorithm are originally formulated for bipartite graphs, but can also be applied to general directed or undirected graphs. Our extensive experimental analysis demonstrates that the proposed algorithm achieves up to 21.26% higher reduction in the number of edges and runs up to 105.18× faster than the FM algorithm. On large synthetic graphs with up to 1.05 billion edges, it attains a reduction in the number of edges of up to 74.36%. On real-world graphs, it achieves a reduction in the number of edges by up to 46.8%. Furthermore, when used as a preprocessing step, our approach yields up to a 2.07× speedup for the matching algorithms on large synthetic graphs, and up to a 1.74× speedup for the All-Pairs Shortest Path algorithms on real-world graphs, when compared to using the given graph as input.
keywords: graph algorithms, speeding-up graph algorithms, clique partitioning
@@ 1 Introduction
Graph algorithms are essential for analyzing complex real-world networks across diverse domains like social interactions, biological systems, and communication infrastructures. However, as these networks grow in size and complexity, the time required to process them using graph algorithms becomes excessively high. This paper tackles the critical issue of high computation time in large-size graphs by introducing a graph restructuring approach that speeds up the execution of graph algorithms such as matching and all pairs shortest paths. This approach involves transforming the graph into a more computationally suitable structure, preserving its path information, where path information refers to the existence of paths between vertices. Specifically, by restructuring the graph to reduce the number of edges, we aim to speeding-up graph algorithms. A key requirement for this is the preservation of complete path information during the graph restructuring process. Many crucial graph algorithms, such as matching and all-pairs shortest paths (APSP), rely on this global connectivity information to produce correct results.
@@
However, to effectively speed up the downstream algorithms, the restructuring algorithm itself must be efficient and the resulting structure readily usable. Specifically, it must meet two critical conditions: 1) the graph restructuring algorithm must have a low running time to ensure that its use as a preprocessing step leads to an overall reduction in execution time compared to running the algorithms on the original graph; and 2) the restructured graph should be directly usable as input to these algorithms or require minimal modifications. These conditions highlight the inherent challenges in designing effective graph restructuring algorithms for speeding-up graph algorithms. In this paper, we address these challenges and propose a new Clique Partitioning based Graph Restructuring (CPGR) algorithm that restructures the graph to reduce the number of edges while preserving the path information. We theoretically and experimentally show that our algorithm achieves faster running time and greater reduction in the number of edges compared to the algorithm proposed by Feder and Motwani [14], i.e., the best existing algorithm used to speeding-up the execution of graph algorithms.
1.1 Related Work
Table 1: Overview of Graph Restructuring Methods
Technique Goal Mechanism Applications Limitations for Speeding-up Graph        Algorithms
Sparsification [15, 18, 20] Reduce # of Edges Edge sub-sampling or filtering Graph        Algorithms,       Visualization Accuracy loss, potential disconnection
Coarsening [6, 7] Reduce # of Vertices Vertex grouping / aggregation Multilevel Methods, GNN Scaling Loss of fine-grained detail, interpolation error
Hierarchical Decomposition [2, 4, 12, 19] Reduce Problem Complexity Subdivide problem using graph structure Optimization, Modeling Applicability depends on problem structure
Lossless Compression [1, 2, 3, 5, 10, 13] Reduce Storage Space Efficient encoding of graph data Storage, Querying Compressed Data Query overhead, less effective on irregular graphs
Clique Partitioning [14], CPGR [this paper] Speeding-up Graph Algorithms Extracting bipartite cliques Graph Algorithms,        Storage
Graph-structured data are common across domains such as social networks, biology, and the web, often reaching billions of edges. To address the computational challenges posed by such scale, graph restructuring techniques have emerged as a key approach for improving the efficiency of graph algorithms through structural transformation. These restructuring methods aim to produce graph representations that are enabling fast computation. These approaches include sparsification, graph coarsening, hierarchical decomposition, and lossless compression, each offering trade-offs between speed and accuracy as shown in Table 1.
Graph sparsification reduces the edge count while approximately preserving properties like cuts, distances, or spectral characteristics, enabling faster execution of algorithms such as PageRank, at the cost of precision [15, 18, 20]. Graph coarsening merges nodes and edges to create a smaller approximation of the original graph, commonly used in multilevel algorithms for partitioning and clustering [6, 7]. While this improves scalability, it can obscure fine-grained structural details. Hierarchical decomposition organizes graphs into nested clusters or layers, facilitating divide-and-conquer approaches in tasks like community detection and multiscale visualization [2, 4, 12, 19]. The main drawback lies in the overhead of maintaining and traversing hierarchical structures. Lossless compression techniques aim to reduce the graph size while preserving exact recoverability, typically using encoding, redundancy elimination, or succinct data structures [1, 2, 3, 5, 10, 13]. Although accurate, such methods can add significant overhead.
Despite their advantages, many of these techniques either compromise path-preservation or introduce substantial preprocessing overhead, limiting their applicability to algorithms that require full connectivity information such as matching and APSP. A more targeted approach is path-preserving graph restructuring, proposed by Feder and Motwani [14], that partitions the graph into bipartite cliques to speeding-up downstream graph algorithms such as bipartite matching, APSP, and vertex connectivity. Their approach guarantees path preservation but suffers from sequential dependencies during clique extraction, which may limit the performance in terms of execution time and reduction in the number of edges.
Building on this foundation, our work proposes a novel Clique Partitioning-based Graph Restructuring (CPGR) algorithm. Unlike Feder and Motwani’s iterative singleton clique extraction, CPGR introduces multiple and larger clique extraction, enabling more efficient restructuring while preserving exact path information. This makes CPGR a promising candidate for preprocessing large-scale graphs to accelerate a wide range of graph algorithms.
@@ 2 Background and Motivation
2.1 Background
Figure 1: (a) Given bipartite graph $G(U,W,E)$ and (b) the tripartite graph that replaces the $\delta$-clique with left partition $\{u_{1},u_{2},u_{3},u_{4},u_{5},u_{7},u_{8}\}$ and right partition $\{w_{4},w_{5}\}$ in the restructured graph $G^{*}(U,W,Z,E^{*})$.
Feder and Motwani [14] proposed a graph restructuring algorithm—referred to as FM in this paper—designed for bipartite graphs $G(U,W,E)$, where $|U|=|W|=n$ and $|E|=m$. The algorithm identifies complete bipartite subgraphs called $\delta$-cliques with left and right partitions of size $\lceil n^{1-\delta}\rceil$ and $k(n,m,\delta)=\big\lfloor\frac{\delta\log n}{\log(2n^{2}/m)}\big\rfloor$, respectively, for some constant $0\leq\delta\leq 1$. Each $\delta$-clique is then replaced by a more compact tripartite graph that introduces a new vertex $z_{q}$, connecting it to all vertices in the clique’s left and right partitions. This transformation reduces the number of edges from $|U_{q}|\times|W_{q}|$ to $|U_{q}|+|W_{q}|$ per clique, achieving significant reduction in the number of edges while preserving the original graph’s connectivity between $U$ and $W$. For instance, Figures 1a and 1b show an example of a $\delta$-clique with left partition $\{u_{1},u_{2},u_{3},u_{4},u_{5},u_{7},u_{8}\}$ and right partition $\{w_{4},w_{5}\}$ in a bipartite graph $G$, and the corresponding tripartite graph that replaces it in the restructured graph $G^{*}$, respectively. The number of edges in the $\delta$-clique is $14=7\times 2$ while the number of edges in the corresponding tripartite graph is $9=7+2$, thus the number of edges in the restructured graph is reduced by 5. Thus, the restructured graph $G^{*}=(U,W,Z,E^{*})$ consists of three disjoint vertex sets, where $Z$ contains one new vertex per extracted clique, and edges in $E^{*}$ connect vertices across these partitions. Although designed for bipartite graphs, the approach can be extended to general graphs as well. Further details on the FM algorithm, including the clique extraction process and the use of neighborhood trees, are provided in Appendix A.1.
2.2 Motivation
We motivate our approach by highlighting three critical aspects related to extracting and replacing $\delta$-cliques that influence both the running time and the reduction in the number of edges. These aspects arise in existing approaches and suggest opportunities for improving how dense substructures are identified and processed.
Vertex selection. A key challenge in identifying $\delta$-cliques for reducing the number of edges lies in selecting vertices for the right partition $W_{q}$ that share a large set of common neighbors in $U$. In approaches such as the FM algorithm, vertices for the right partition are selected using a combinatorial function that accounts for vertex degrees across different levels of a neighborhood tree. At each level of the neighborhood tree, vertex degrees within the corresponding partitions are aggregated to determine which branch to explore. As a result, vertex selection is driven by combined degree contributions across partitions rather than by the individual potential of vertices to form dense bipartite subgraphs with $U$, which may cause vertices with higher degree to be overlooked when forming dense $\delta$-cliques.
Figure 2: Progression of $\hat{m}$, $\hat{k}$, and number of cliques extracted for a graph with 128 vertices in each bi-partition, density 0.98, and $\delta=1$ by (a) FM and (b) CPGR.
Extracting $\delta$-cliques. The process of forming $\delta$-cliques relies on selecting $\hat{k}=k(n,\hat{m},\delta)$ vertices from the right partition $W$, where $\hat{m}$ is the number of remaining edges in the graph and $n$ is the number of vertices in each partition. As the algorithm progresses and more cliques are extracted, the number of edges $\hat{m}$ decreases, which in turn affects the value of $\hat{k}$. Due to the stepwise nature of the function $k(n,\hat{m},\delta)$, the value of $\hat{k}$ often remains constant across multiple iterations. As illustrated in Figure 2a, this results in a staircase-like progression in which several iterations operate under the same $\hat{k}$ while removing relatively few edges per iteration. This behavior suggests that the iterative structure may not fully exploit the potential of each $\hat{k}$ value, particularly during phases in which $\hat{k}$ remains unchanged over multiple iterations.
Update mechanism. After selecting vertices from the right partition, the manner and timing of degree updates play a crucial role in determining which edges remain available for future $\delta$-clique discovery. In the existing approach (FM algorithm), degree updates are applied incrementally as candidate vertices are selected, rather than after a $\delta$-clique has been fully determined. As a result, degree reductions may be applied to vertices based on partial clique construction, rather than solely on edges that are ultimately removed by an extracted clique. Such early updates can reduce the availability of edges that could participate in subsequent $\delta$-cliques, particularly in sparse or moderately dense graphs. This observation highlights the importance of update mechanisms that accurately reflect only the edges eliminated by an extracted clique.
2.3 Our contributions
This paper introduces a novel algorithm, the Clique Partitioning-based Graph Restructuring (CPGR) algorithm. Our main contributions are summarized as follows:
  1. 1.
    We present CPGR, a graph restructuring algorithm that removes multiple $\delta$-cliques within a single iteration, achieving a running time of $O(mn^{\delta})$. This improves upon the $O(mn^{\delta}\log^{2}n)$ running time of the FM algorithm.
  2. 2.
    CPGR allows vertices to participate in multiple clique extractions during the restructuring process. For a fixed $\hat{k}$, the algorithm is designed to extract as many valid $\delta$ cliques as possible in each iteration, which increases the total number of extracted cliques, particularly in large and high density graphs, and leads to a larger number of edges removed per iteration, as illustrated in Figure 2b.
  3. 3.
    CPGR achieves an average reduction in the number of edges that is at least as good as the FM algorithm.
  4. 4.
    Experimental analysis show that CPGR achieves up to 21% higher reduction in the number of edges and runs up to 105.18× faster than the FM algorithm.
  5. 5.
    Experimentally, we show the practical utility of CPGR as a preprocessing step, achieving a speedup of up to 2.07× for matching and 1.74× for APSP on different graph datasets.
3 The Proposed Algorithm
3.1 Clique Partitioning-based Graph Restructuring (CPGR) Algorithm
This section details the design of our Clique Partitioning-based Graph Restructuring (CPGR) algorithm, presented in Algorithm 1. For a given input graph $G$, CPGR obtains a restructured graph $G^{*}$ of $G$ by iteratively finding bipartite cliques (i.e., complete bipartite subgraphs) in $G(U,W,E)$ and replacing them with corresponding tripartite graphs until finding new bipartite cliques does not contribute to the reduction in the number of edges of $G$. The size of the right partition of the cliques is determined by $\hat{k}=k(n,\hat{m},\delta)$, which guarantees a $\delta$-clique, and the selection of vertices $w_{j}\in W$ depends on the degree of vertices $d_{w_{j}}=|N(w_{j})|$, where $N(w_{j})$ denotes the set of neighbors of vertex $w_{j}$. The partition of $G$ into bipartite cliques guarantees that each edge in $G$ is in exactly one clique, i.e., $E(C_{i})\cap E(C_{j})=\emptyset$, $\forall i,j$ and $i\neq j$, where $E(C_{i})$ is the set of edges of $\delta$-clique $C_{i}$. The input of CPGR consists of the adjacency matrix $A$ of $G$, and a constant $\delta$.
Initialization (Lines 1-5). For the given bipartite graph $G(U,W,E)$, CPGR initializes the index $q$ of the bipartite cliques extracted from $G$, the number of vertices, $n$ where $n=|U|=|W|$, the number of edges $\hat{m}$, and the degree of the vertices $d_{w_{j}}$} $\forall j=1,\ldots,|W|$, initialized to $0]_{n}$, the zero vector fo size $n$. CPGR also initializes the set of extracted bipartite cliques $\mathcal{C}$, to the empty set.
Algorithm 1 CPGR: Clique Partitioning-based Graph Restructuring Algorithm
1:AA: Adjacency matrix;
2:  $\delta$: a constant such that $0\leq\delta\leq 1$.